OpenStudy (anonymous):

improper integral problem:

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{-\infty}1/(3\sqrt{x})\]

OpenStudy (amistre64):

I spose we might have to break this apart into 2 improps?

OpenStudy (anonymous):

So far I integrate to get: \[2\sqrt{x}/3 \int\limits_{-\infty}^{-\infty}\]

OpenStudy (amistre64):

is this really from -inf to -inf?

OpenStudy (anonymous):

Yea

OpenStudy (amistre64):

any integral from a to a = 0 right?

OpenStudy (anonymous):

So i am not sure how I would evaluate since its negative going into a square root.

OpenStudy (anonymous):

Hmm, not sure

OpenStudy (amistre64):

\[\int_{5}^{5}e^{x^2}dx = 0 \]

OpenStudy (anonymous):

there is no such thing as \[\int_{-\infty}^{-\infty}\]

OpenStudy (amistre64):

\[\int_{124}^{124}x^2+sin(3x^2+4)dx=0\]

OpenStudy (anonymous):

yes but infinity is not a number

OpenStudy (anonymous):

What I am supposed to do is determine whether the problem "converges" or "diverges"

OpenStudy (amistre64):

well, as "b" moves to infinity the limit stuff implies that its 0 or something right?

OpenStudy (anonymous):

i mean maybe you are supposed to say 0, but it doesn't make any sense. an improper integral is a limit

OpenStudy (anonymous):

I don't know thats what I need to find out

OpenStudy (anonymous):

whether the limit approaches infinity (diverges) or a number (converges)

OpenStudy (anonymous):

well i don't know where this question comes from, but if you are really supposed to give an answer just say 0 and don't give it another minutes thought. but if you asked me i would say it does not exist

OpenStudy (amistre64):

\[\int_{b}^{b}f(x)dx=0\] as b appraoches infinity, the integral = 0 still right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So suppose the bottom bound was 0 instead of -infinity

OpenStudy (anonymous):

Then I would have to do the methodology I was thinking of, since we couldn't just say 0 by that theorem.

OpenStudy (amistre64):

then youd be evaluating a different integral :)

OpenStudy (anonymous):

So how could I determine the value limit if when plugging in negative infinity yields an imaginary numer?

OpenStudy (amistre64):

\[\int_{b}^{0}f(x)dx=N\] \[\int_{0}^{b}f(x)dx=-N\] add them together to get 0 id assume

OpenStudy (amistre64):

i havent delved into imaginaries yet :)

OpenStudy (anonymous):

Ah, ok

OpenStudy (anonymous):

I am thinking mayeb l'hopital's rule

OpenStudy (amistre64):

perhaps, but I cant be certain of it .... yet